# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v4

@@ 4.2 Feature space
Synthesizing in raw pixel space often overburdens models by requiring them to reconstruct images from scratch, a task particularly challenging for high-resolution video datasets. This realization has led to a shift in focus among researchers. Rather than grappling with the complexities of pixel-level synthesis, several studies have pivoted towards high-level feature synthesis in feature space, such as segmentation and depth maps. These approaches offer a more efficient way of handling the intricacies of videos [76].
@@
Future semantic segmentation represents a progressive approach of video synthesis, primarily focusing on synthesizing semantic maps for forthcoming video frames. This methodology diverges from traditional raw pixel forecasting, turning to semantic maps to narrow the synthesis scope and enrich scene comprehension. In this context, the S2S model [72] stands as a groundbreaking end-to-end system. It processes RGB frames alongside their semantic maps, both as input and output. This integration not only advances future semantic segmentation but also elevates the task of video frame prediction, showcasing the distinct advantage of semantic-level forecasting. Building on this foundation, SADM [7] further innovates by amalgamating optical flow with semantic maps. This fusion leverages optical flow for motion tracking and semantic maps for appearance detailing, employing the former to warp input images and the latter to inpaint occluded areas.
@@
Depth maps, as a 2D data structure containing 3D information, can provide the model with an enhanced perception of the 3D world at minimal cost. Leveraging predictions of future depth maps can assist in future frame synthesis tasks. MAL [68] introduces a meta-learning framework with a two-branch architecture containing future depth prediction and an auxiliary task of image reconstruction. The proposed meta-learning framework improves the quality of synthesized future frames, especially in complex dynamic scenes.
@@ 5 Stochastic synthesis
In the early stages, video synthesis is primarily perceived as a low-level computer vision task, with a focus on employing deterministic algorithms to enhance pixel-level metrics such as MSE, PSNR, and SSIM. This approach, however, inherently limits the potential for creative output from these models by confining possible motion outcomes to a single, fixed result [76]. This often results in high pixel-level scores but at the cost of producing blurry images, which significantly diminishes the practical applicability of these algorithms. Recognizing this issue, there has been a paradigm shift in the field of video synthesis, moving away from the reliance on short-term deterministic prediction towards embracing long-term stochastic generation. This transition acknowledges that while stochastic synthesis may yield results that diverge significantly from the ground truth, it plays a crucial role in fostering a more comprehensive understanding and enhancing creativity concerning the evolution of video content.
@@ 5.1.1 Stochastic distributions
In the early stage, VPN [56] uses CNNs for multiple predictions in videos based on pixel distributions and SV2P [4] enhances an action-conditioned model [31] with stochastic distribution estimation for videos. Shifting the focus to a more holistic view of video elements, PFP model [48] presents a probabilistic method for synthesizing semantic segmentation, depth map, and optical flow in videos simultaneously. Additionally, SRVP [33] utilizes Ordinary Differential Equations (ODEs), while PhyDNet [41] employs Partial Differential Equations (PDEs) to calculate stochastic distributions.
5.1.2 Probabilistic models
@@
With the pioneering work of GDL [74] establishing the foundation, adversarial training has significantly advanced future frame synthesis task in predicting uncertain object motions. Similarly, vRNN [16] and GHVAE [114] enhance VAEs through likelihood networks and hierarchical structures respectively, thereby offering another dimension to the ongoing evolution of stochastic synthesis methodologies.
@@
Recognizing that object motion is largely deterministic barring unforeseen events like collisions, SVG [24] models trajectory uncertainty using fixed and learnable priors, which effectively blends deterministic and probabilistic approaches. In a similar vein, but with a focus on enhancing temporal aspects, Retrospective Cycle GAN [62] introduces a sequence discriminator designed to detect fake frames. This idea of scrutinizing frame authenticity is further extended in DIGAN [129], where the emphasis shifts to a motion discriminator that concentrates on identifying unnatural motions.
@@
To overcome pixel-level synthesis challenges in stochastic models, several works introduce intermediate representations. S2S [72] and Vid2Vid [102] integrate adversarial training with future semantic segmentation. Additionally, TPK model [101] leverages a VAE to extract human pose information, and then a GAN for predicting future poses and frames.
@@
It is worth noting that directly modeling stochastic distributions is prone to covering a broader predictive distribution but suffers from poor visual effects. In contrast, probabilistic models are capable of producing sharper results, yet they grapple with issues such as mode collapsing, training difficulties, and substantial computational overhead. Bridging these two methodologies, SAVP [63] combines stochastic distributions with adversarial training aiming to achieve broader predictive distributions with promising visual quality.
@@ 5.2.1 Content and motion
Video synthesis algorithms confront the challenging complexity of natural video sequences by emphasizing intricate image details. To this end, they focus on effectively modeling appearances through detailed local information, while simultaneously developing a comprehensive understanding of the dynamic global content inherent in videos. However, in applications such as robot navigation and autonomous driving, the importance of understanding object motion patterns supersedes the pursuit of visual aesthetics. This priority shift has encouraged the development of algorithms that emphasize predicting the movement of objects and distinguishing motion from appearance in videos. Early work CDNA [31] has set a precedent by explicitly predicting the motion of objects. It maintains invariant appearance characteristics, which is instrumental in applying the model to objects beyond those encountered during training.
@@
MoCoGAN [95] automatically learns to disentangle motion from content in an unsupervised manner, while the approach of leveraging separate content and motion encoder pathways has also been widely used in various video prediction models. This idea of separating content and motion is further explored in LMC [64], which makes the motion encoder focused on motion prediction based on residual frames, while the content encoder extracts content features from the input frame sequence. MMVP [133] takes a different approach by employing only one image encoder to extract information, and then a two-stream network before the image decoder to separately handle motion prediction and appearance maintenance.
@@
Addressing the stochastic nature of motion, AMC-GAN [53] models multiple plausible outcomes via adversarial training. Transitioning to a different approach, SLAMP [3] adopts a non-adversarial approach but focuses on learning stochastic variables for separate content and motion. Further advancing this field, LEO [111] and D-VDM [84] leverage diffusion models for more realistic disentanglement of content and motion, showcasing the latest advancements in this direction.
5.2.2 Foreground and background
@@
In the process of predicting future frames, the motion dynamics of the foreground (objects) and background (scene) typically exhibit significant differences. Foreground objects typically exhibit more intense movement, whereas scenes tend to remain relatively static. This distinction has steered research towards predicting the motions of these elements separately, offering a more nuanced understanding of video dynamics. A pivotal contribution in this area is DrNet [25], which specifically addresses scenarios where the background remains largely unchanged across video frames. This model ingeniously decomposes images into the content and pose of objects, then harnesses adversarial training techniques to develop a scene discriminator that evaluates whether two pose vectors belong to the same video sequence. Similarly, OCVP-VP [98] utilizes a slot-wise scene parsing network SAVi [58] to achieve segmentation from the scene to the object level. Focusing on such videos, prediction models can streamline the prediction process by eliminating the need to learn complex scene dynamics. Human-centric tasks, such as predicting human movement and interaction with the environment, and object-centric tasks, such as tracking object motion and positioning, both benefit from this approach.
@@
Future frame synthesis often centers around foreground motion, particularly when it involves intricate human movements. In such scenarios, a common assumption across various specialized datasets is the relatively static nature of the background, which is a typical characteristic in datasets focusing on detailed human motion. This has led to a significant focus in research on understanding and forecasting human poses to enhance the prediction of foreground motion. An example of this approach is seen in the work of DVGPC [13], which innovatively predicts skeleton motion sequences before transforming these sequences into pixel space using a skeleton-to-image transformer. Their method is an effective solution for linking the abstract representation of motion with the practical aspect of video prediction, demonstrating a nuanced understanding of the complexities involved in human-centric FFS tasks.
@@
The field of object-centric video prediction is first introduced in the work of CVP [124], which lays the foundation for this specialized area of video prediction. SlotFormer [118] introduces transformer-based autoregressive models to learn representations of each object in video sequences. This innovation ensures consistent and accurate tracking of each object over time. The OKID model [21] represents a recent advancement. It uniquely decomposes videos into distinct elements, specifically the attributes and trajectory dynamics of moving objects, employing a Koopman operator. This approach highlights a more detailed method of analyzing object motion within video sequences, differentiating it from previous methodologies.
@@
Methods that concentrate on human poses or objects have demonstrated considerable promise in specific video datasets, yet they encounter limitations due to their reliance on pre-defined structures and struggle with adapting to variable backgrounds, which hampers their ability to generalize. This challenge is evident in their performance, which, while effective under certain conditions, falters when faced with dynamic background shifts, revealing a lack of the versatility required for wider applications. Bridging this gap, MOSO [92] emerges as a notable approach, identifying motion, scene, and object as the pivotal elements of a video. It delves deeper into content analysis by distinguishing between scene and object. Scene and object are considered as a further breakdown of content, where scene represents the background and object represents the foreground. MOSO’s innovative contribution lies in its two-stage network tailored for general video analysis. Initially, the MOSO-VQVAE model dissects video frames into token-level representations, honing its capabilities through a video reconstruction task. Subsequently, the model employs transformers in its second stage, tackling disparities in masked tokens. This strategic design equips the model to handle a variety of tasks at the token level, including video prediction, interpolation, and unconditional video generation.
@@ 5.3 Motion-controllable synthesis
In the field of future frame synthesis, one specialized research direction has emerged that focuses on the explicit control of motion. This approach is distinct in its emphasis on forecasting future object positions based on user-defined instructions, diverging from the conventional reliance on past motion trends. The central challenge in this domain lies in the synthesis of videos that adhere to these direct instructions while maintaining a natural and coherent flow, a task that demands a nuanced understanding of both user intent and the dynamics of motion within a video context. This challenge highlights the intricate balance between user control and automated imagination, marking a significant shift in how future frame synthesis models are conceptualized and implemented.
@@
There is no historical motion information to be used for video generation from one still image, hence several methods have emerged allowing for interactive user control. iPOKE [8] introduces techniques where local interactive strokes and pokes enable users to deform objects in one still image to generate a sequence of video frames. These strokes indicate the user’s intended motion for the objects. Following this innovative pathway, Controllable-Cinemagraphs model [73] introduces a method to interactively control the animation of fluid elements. The advancements in the field underscore the growing importance of user-centric approaches in the realm of motion-controllable future frame synthesis.
@@
The integration of instructions of various modalities, including local strokes, sketches, and texts, is increasingly common in works aiming to capture user-specified motion trends. VideoComposer [104] synthesizes videos by combining text descriptions, hand-drawn strokes, and sketches. This approach respects textual, spatial, and temporal constraints, leveraging video latent diffusion models and motion vectors for explicit dynamic guidance. Essentially, It can generate videos that align with user-defined motion strokes and shape sketches. In a similar vein, DragNUWA [125] primarily leverages text for video content description and strokes for future motion control, enabling the generation of customizable videos. These attempts advance the field of video generation by broadening the range of user input modalities.

### v5

@@ 3.2 Feature Space
Synthesizing in raw pixel space often overburdens models by requiring them to reconstruct images from scratch, a task particularly challenging for high-resolution video datasets. This realization has led to a shift in focus among researchers. Rather than grappling with the complexities of pixel-level synthesis, several studies have pivoted towards high-level feature synthesis in feature space, such as segmentation and depth maps. These approaches offer a more efficient way of handling the intricacies of videos (Oprea et al., 2020).
@@
Future semantic segmentation represents a progressive approach of video synthesis, primarily focusing on synthesizing semantic maps for forthcoming video frames. This methodology diverges from traditional raw pixel forecasting, turning to semantic maps to narrow the synthesis scope and enrich scene comprehension. In this context, the S2S model (Luc et al., 2017) stands as a groundbreaking end-to-end system. It processes RGB frames alongside their semantic maps, both as input and output. This integration not only advances future semantic segmentation but also elevates the task of video frame prediction, showcasing the distinct advantage of semantic-level forecasting. Building on this foundation, SADM (Bei et al., 2021) further innovates by amalgamating optical flow with semantic maps. This fusion leverages optical flow for motion tracking and semantic maps for appearance detailing, employing the former to warp input images and the latter to inpaint occluded areas.
@@
Depth maps, as a 2D data structure containing 3D information, can provide the model with an enhanced perception of the 3D world at minimal cost. Leveraging predictions of future depth maps can assist in FFS tasks. MAL (Liu et al., 2023) introduces a meta-learning framework with a two-branch architecture containing future depth prediction and an auxiliary task of image reconstruction. The proposed meta-learning framework improves the quality of synthesized future frames, especially in complex dynamic scenes.
Challenges.
Future prediction in feature space, presents significant challenges due to the complex interplay of temporal dynamics and spatial context. Models must capture intricate motion patterns and predict accurate depths or semantic regions, requiring a deep understanding of 3D scene structures and object interactions. Ensuring temporal consistency and precise spatial accuracy while handling occlusions, perspective changes, and complex backgrounds is crucial. High-resolution feature maps and large-scale annotated datasets further increase computational demands and data requirements. Generalizing to unseen scenes and objects remains a formidable challenge, necessitating robust models that can adapt to diverse visual appearances and contexts. These challenges underscore the need for innovative approaches like meta-auxiliary learning to enhance future prediction capabilities.
@@ 4 Stochastic Synthesis
In the early stages, video synthesis is primarily perceived as a low-level computer vision task, with a focus on employing deterministic algorithms to enhance pixel-level metrics such as MSE, PSNR, and SSIM. This approach, however, inherently limits the potential for creative output from these models by confining possible motion outcomes to a single, fixed result (Oprea et al., 2020). This often results in high pixel-level scores but at the cost of producing blurry images, which significantly diminishes the practical applicability of these algorithms. Recognizing this issue, there has been a paradigm shift in the field of video synthesis, moving away from the reliance on short-term deterministic prediction towards embracing long-term stochastic generation. This transition acknowledges that while stochastic synthesis may yield results that diverge significantly from the ground truth, it plays a crucial role in fostering a more comprehensive understanding and enhancing creativity concerning the evolution of video content.
@@ 4.1 Stochasticity Modelling
In the early stage, VPN (Kalchbrenner et al., 2017) uses CNNs for multiple predictions in videos based on pixel distributions and SV2P (Babaeizadeh et al., 2017) enhances an action-conditioned model (Finn et al., 2016) with stochastic distribution estimation for videos. Shifting the focus to a more holistic view of video elements, PFP model (Hu et al., 2020) presents a probabilistic method for synthesizing semantic segmentation, depth map, and optical flow in videos simultaneously. Additionally, SRVP (Franceschi et al., 2020) utilizes Ordinary Differential Equations (ODEs), while PhyDNet (Guen & Thome, 2020) employs Partial Differential Equations (PDEs) to calculate stochastic distributions. A potential drawback is that their assumption that physical laws can be linearly disentangled from other factors of variation in the latent space, may not hold true for all types of videos.
Probabilistic models.
@@
With the pioneering work of GDL (Mathieu et al., 2015) establishing the foundation, adversarial training has significantly advanced FFS tasks in predicting uncertain object motions. Similarly, vRNN (Castrejon et al., 2019) and GHVAE (Wu et al., 2021) enhance VAEs through likelihood networks and hierarchical structures respectively, thereby offering another dimension to the ongoing evolution of stochastic synthesis methodologies.
@@
Recognizing that object motion is largely deterministic barring unforeseen events like collisions, SVG (Denton & Fergus, 2018) models trajectory uncertainty using fixed and learnable priors, which effectively blends deterministic and probabilistic approaches. In a similar vein, but with a focus on enhancing temporal aspects, Retrospective Cycle GAN (Kwon & Park, 2019) introduces a sequence discriminator designed to detect fake frames. This idea of scrutinizing frame authenticity is further extended in DIGAN (Yu et al., 2022), where the emphasis shifts to a motion discriminator that concentrates on identifying unnatural motions.
@@
To overcome pixel-level synthesis challenges in stochastic models, several works introduce intermediate representations. S2S (Luc et al., 2017) and Vid2Vid (Wang et al., 2018a) integrate adversarial training with future semantic segmentation. Additionally, TPK model (Walker et al., 2017) leverages a VAE to extract human pose information, and then a GAN for predicting future poses and frames.
@@
It is worth noting that directly modeling stochastic distributions is prone to covering a broader predictive distribution but suffers from poor visual effects. In contrast, probabilistic models are capable of producing sharper results, yet they grapple with issues such as mode collapsing, training difficulties, and substantial computational overhead. Bridging these two methodologies, SAVP (Lee et al., 2018) combines stochastic distributions with adversarial training aiming to achieve broader predictive distributions with promising visual quality.
Challenges.
Despite the ability of stochastic models to capture a broad spectrum of plausible futures, they often grapple with visual quality issues and heightened computational demands. Directly modeling stochastic distributions can result in blurred outputs, while probabilistic models may face challenges such as mode collapse and training instability. The quest for a harmonious balance between diverse, high-quality predictions and computational efficiency remains a formidable challenge. Moreover, the assumption that physical laws can be linearly disentangled from other variation factors may not universally apply to all video types, hinting at the need for more adaptable and generalizable models in the future.
@@ 4.2 Disentangling Components
Video synthesis algorithms confront the challenging complexity of natural video sequences by emphasizing intricate image details. To this end, they focus on effectively modeling appearances through detailed local information, while simultaneously developing a comprehensive understanding of the dynamic global content inherent in videos. However, in applications such as robot navigation and autonomous driving, the importance of understanding object motion patterns supersedes the pursuit of visual aesthetics. This priority shift has encouraged the development of algorithms that emphasize predicting the movement of objects and distinguishing motion from appearance in videos. Early work CDNA (Finn et al., 2016) has set a precedent by explicitly predicting the motion of objects. It maintains invariant appearance characteristics, which is instrumental in applying the model to objects beyond those encountered during training.
@@
MoCoGAN (Tulyakov et al., 2018) automatically learns to disentangle motion from content in an unsupervised manner, while the approach of leveraging separate content and motion encoder pathways has also been widely used in various video prediction models. This idea of separating content and motion is further explored in LMC (Lee et al., 2021), which makes the motion encoder focused on motion prediction based on residual frames, while the content encoder extracts content features from the input frame sequence. MMVP (Zhong et al., 2023) takes a different approach by employing only one image encoder to extract information, and then a two-stream network before the image decoder to separately handle motion prediction and appearance maintenance.
@@
Addressing the stochastic nature of motion, AMC-GAN (Jang et al., 2018) models multiple plausible outcomes via adversarial training. Transitioning to a different approach, SLAMP (Akan et al., 2021) adopts a non-adversarial approach but focuses on learning stochastic variables for separate content and motion. Further advancing this field, LEO (Wang et al., 2023e) and D-VDM (Shen et al., 2023) leverage diffusion models for more realistic disentanglement of content and motion, showcasing the latest advancements in this direction.
Foreground and background.
@@
In the process of predicting future frames, the motion dynamics of the foreground (objects) and background (scene) typically exhibit significant differences. Foreground objects typically exhibit more intense movement, whereas scenes tend to remain relatively static. This distinction has steered research towards predicting the motions of these elements separately, offering a more nuanced understanding of video dynamics. A pivotal contribution in this area is DrNet (Denton et al., 2017), which specifically addresses scenarios where the background remains largely unchanged across video frames. This model ingeniously decomposes images into the content and pose of objects, then harnesses adversarial training techniques to develop a scene discriminator that evaluates whether two pose vectors belong to the same video sequence. Similarly, OCVP-VP (Villar-Corrales et al., 2023) utilizes a slot-wise scene parsing network SAVi (Kipf et al., 2021) to achieve segmentation from the scene to the object level. Focusing on such videos, prediction models can streamline the prediction process by eliminating the need to learn complex scene dynamics. Human-centric tasks, such as predicting human movement and interaction with the environment, and object-centric tasks, such as tracking object motion and positioning, both benefit from this approach.
@@
FFS often centers around foreground motion, particularly when it involves intricate human movements. In such scenarios, a common assumption across various specialized datasets is the relatively static nature of the background, which is a typical characteristic in datasets focusing on detailed human motion. This has led to a significant focus in research on understanding and forecasting human poses to enhance the prediction of foreground motion. An example of this approach is seen in the work of DVGPC (Cai et al., 2018), which innovatively predicts skeleton motion sequences before transforming these sequences into pixel space using a skeleton-to-image transformer. Their method is an effective solution for linking the abstract representation of motion with the practical aspect of video prediction, demonstrating a nuanced understanding of the complexities involved in human-centric FFS tasks.
@@
The field of object-centric video prediction is first introduced in the work of CVP (Ye et al., 2019), which lays the foundation for this specialized area of video prediction. SlotFormer (Wu et al., 2022c) introduces transformer-based auto-regressive models to learn representations of each object in video sequences. This innovation ensures consistent and accurate tracking of each object over time. The OKID model (Comas et al., 2023) represents a recent advancement. It uniquely decomposes videos into distinct elements, specifically the attributes and trajectory dynamics of moving objects, employing a Koopman operator. This approach highlights a more detailed method of analyzing object motion within video sequences, differentiating it from previous methodologies.
@@
Methods that concentrate on human poses or objects have demonstrated considerable promise in specific video datasets, yet they encounter limitations due to their reliance on pre-defined structures and struggle with adapting to variable backgrounds, which hampers their ability to generalize. This challenge is evident in their performance, which, while effective under certain conditions, falters when faced with dynamic background shifts, revealing a lack of the versatility required for wider applications. Bridging this gap, MOSO (Sun et al., 2023) emerges as a notable approach, identifying motion, scene, and object as the pivotal elements of a video. It delves deeper into content analysis by distinguishing between scene and object. Scene and object are considered as a further breakdown of content, where scene represents the background and object represents the foreground. MOSO’s innovative contribution lies in its two-stage network tailored for general video analysis. Initially, the MOSO-VQVAE model dissects video frames into token-level representations, honing its capabilities through a video reconstruction task. Subsequently, the model employs transformers in its second stage, tackling disparities in masked tokens. This strategic design equips the model to handle a variety of tasks at the token level, including video prediction, interpolation, and unconditional video generation.
Challenges.
Disentangling content from motion or foreground from background in videos is complex due to the intricate interplay between temporal dynamics and spatial context. Models must accurately capture and predict motion patterns, depths, and semantic regions while maintaining temporal consistency and spatial accuracy. Handling occlusions, perspective changes, and complex backgrounds adds to the difficulty. High-resolution feature maps and large-scale annotated datasets exacerbate computational demands. Generalizing to unseen scenes and objects remains challenging, necessitating robust models that can adapt to diverse visual appearances and contexts. Applying the concept of separate processing to the generative methods we will discuss later (in Section 5) is also a challenge.
@@ 4.3 Motion-Controllable Synthesis
In the field of FFS, one specialized research direction has emerged that focuses on the explicit control of motion. This approach is distinct in its emphasis on forecasting future object positions based on user-defined instructions, diverging from the conventional reliance on past motion trends. The central challenge in this domain lies in the synthesis of videos that adhere to these direct instructions while maintaining a natural and coherent flow, a task that demands a nuanced understanding of both user intent and the dynamics of motion within a video context. This challenge highlights the intricate balance between user control and automated imagination, marking a significant shift in how FFS models are conceptualized and implemented.
@@
There is no historical motion information to be used for video generation from one still image, hence several methods have emerged allowing for interactive user control. iPOKE (Blattmann et al., 2021) introduces techniques where local interactive strokes and pokes enable users to deform objects in one still image to generate a sequence of video frames. These strokes indicate the user’s intended motion for the objects. Following this innovative pathway, Controllable-Cinemagraphs model (Mahapatra & Kulkarni, 2022) introduces a method to interactively control the animation of fluid elements. The advancements in the field underscore the growing importance of user-centric approaches in the realm of motion-controllable FFS.
@@
The integration of instructions of various modalities, including local strokes, sketches, and texts, is increasingly common in works aiming to capture user-specified motion trends. VideoComposer (Wang et al., 2023a) synthesizes videos by combining text descriptions, hand-drawn strokes, and sketches. This approach respects textual, spatial, and temporal constraints, leveraging video latent diffusion models and motion vectors for explicit dynamic guidance. Essentially, It can generate videos that align with user-defined motion strokes and shape sketches. In a similar vein, DragNUWA (Yin et al., 2023) primarily leverages text for video content description and strokes for future motion control, enabling the generation of customizable videos. These attempts advance the field of video generation by broadening the range of user input modalities.
Challenges.
Achieving natural and coherent video synthesis under explicit user control is challenging. Models must interpret user intent accurately and generate videos that adhere to specified motion instructions while maintaining temporal and spatial coherence. Balancing user control with the model’s autonomous imagination is crucial. Ensuring that the generated videos are both visually appealing and contextually appropriate adds to the complexity, requiring a deep understanding of both user input and video dynamics.